# Boundary Constraint-free Biomechanical Model-Based Surface Matching for Intraoperative Liver Deformation Correction

### v3

@@ Abstract
In image-guided liver surgery, 3D-3D non-rigid registration methods play a crucial role in estimating the mapping between the preoperative model and the intraoperative surface represented as point clouds, addressing the challenge of tissue deformation. Typically, these methods incorporate a biomechanical model, represented as a finite element model (FEM), used to regularize a surface matching term. This paper introduces a unique 3D-3D non-rigid registration method, which, in contrast to the preceding techniques. Our method incorporates the FEM within the surface-matching term itself, inherently ensuring that the estimated deformation maintains geometric consistency throughout the registration process. Additionally, we eliminate the need to identify zero-boundary conditions and applied force locations in the FEM. We achieve this by integrating soft springs into the stiffness matrix and allowing forces to be distributed across the entire liver surface. Optimization is achieved through an accelerated gradient algorithm, further enhanced by our proposed method for determining the optimal step size. We evaluated our method and compared it to several state-of-the-art methods across various datasets. Our straightforward and effective approach consistently outperformed or achieved comparable performance to the state-of-the-art methods. Our code and datasets are available at https://github.com/zixinyang9109/BCF-FEM.
@@ I-B Contributions
Most non-rigid registration methods incorporate the FEM as the strain energy/regularization term. Only few methods include the FEM as the surface matching data term, which requires the boundary conditions to be manually identified. To mitigate the inconvenience of the manual intervention, which is also prone to error, our proposed method incorporates the modified FEM into the data term using a modified stiffness matrix. This approach stabilizes the stiffness matrix and eliminates the need for manual identification of ZBCs and fore locations, hence rendering our method unique among the existing state-of-the-art methods.
@@ II-A Biomechanical Model
is a matrix used in each iteration of the solution of a nonlinear problem, and it changes with each iteration. For a linear elastic material, the tangent stiffness matrix is simply the stiffness matrix.
@@ III Methods
Our method employs a force-driven approach to deform the source model for optimal alignment with the target point cloud. Fig. 1 provides an overview of our registration method. The objective function comprises a surface matching term. The surface matching term utilizes a correspondence matrix generated by a closest point operator and a FEM. Minimization of the objective function is achieved through an accelerated gradient algorithm. During the force optimization phase, we initially determine the optimal step for calculating Nesterov’s accelerated gradient [29]. Subsequently, the gradient is employed to update forces. In each iteration, we first update the correspondence matrix via the closest point operator, then update the force through force optimization; finally, we update the displacements using the FEM. After a set number of iterations, the predicted volumetric deformation enables the preoperative volume model to align with the intraoperative point cloud. In the following section, we will detail each of these steps.
@@ III-A Objective Function
where $\mathbf{C}\in\mathbb{R}^{3m\times 3n}$ is a soft correspondence matrix, $\mathbf{x}\in\mathbb{R}^{3n}$ are spatial location of the preoperative mesh nodes, $\mathbf{u}\in\mathbb{R}^{3n}$ are displacements associated with the nodes, and $\mathbf{y}\in\mathbb{R}^{3m}$ is the intraoperative point cloud. The correspondence matrix $\mathbf{C}$ maps the deformed to a point on the surface will be detailed in Sec. III-B. Our data term is a reformulation of the data term from Mestdagh et al. [16, 19] where we have explicitly incorporated the correspondence matrix and the dependence of $J(\mathbf{u})$ on $\mathbf{u}$.
@@ III-C Optimization
Nesterov’s accelerated gradient is given by Eq. 6 and Eq. 7. $\mathbf{p}^{k}$ carries momentum from the previous iteration by combining the force estimates of the current and previous iterations. Eq. 7 updates the forces, where $\nabla J(\mathbf{p}^{k})$ is the gradient of the data term (Eq. 4) with respect to $\mathbf{p}$, and $\alpha^{k}$ is the step size. Using the relationship between $\mathbf{u}$ and $\mathbf{f}$ given by Eq.2, the $\nabla J(\mathbf{p}^{k})$ is given by
@@
Setting the derivative of Eq. 10 for $\alpha$ to zero, we can obtain the optimal step size
@@ III-D Modifications for Unknown Zero Boundary Conditions and Force Locations
where $k_{ss}$ is the stiffness of the soft spring and $\mathbf{I}$ is the identity matrix. This is analogous to the work presented by Peterlik et al. [3], where a damping term was added to the stiffness matrix, which was then integrated into a strain/deformation energy term to regularize a data term. In our work, the damping term is integrated directly into the date term via the FEM. The added small stiffness increases the inertia of the liver model being deformed, preventing rigid motion without internal deformations. This approach enables us to allow the forces to be applied at any spatial location on the liver surface.
@@ IV-2 In vitro phantoms
Additionally, partial anterior surfaces were cropped to achieve surface visibility ratios ranging from approximately $20\%$ to $27\%$, calculated as the ratios of cropped surface areas to the full surface areas, as summarized in Table I.
@@ IV-3 Open-CAS dataset
: The Open-CAS dataset contains three in silico phantoms and one in vitro phantom, as described in detailed in [5]. Deformations of the in silico phantoms are generated using a non-linear biomechanical model. The in vitro phantom featuring 6 fiducial markers was fabricated from silicone, and two intraoperative surfaces from a CT and stereo-endoscopic imaging are included.

### v4

@@ Abstract
In image-guided liver surgery, 3D-3D non-rigid registration methods play a crucial role in estimating the mapping between the preoperative model and the intraoperative surface represented as point clouds, addressing the challenge of tissue deformation. Typically, these methods incorporate a biomechanical model, represented as a finite element model (FEM), into the strain energy term to regularize a surface matching term. We propose a 3D-3D non-rigid registration method that incorporates a modified FEM into the surface matching term. The modified FEM alleviates the need to specify boundary conditions, which is achieved by modifying the stiffness matrix of a FEM and using diagonal loading for stabilization. As a result, the modified surface matching term does not require the specification of boundary conditions or an additional strain energy term to regularize the surface matching term. Optimization is achieved through an accelerated gradient algorithm, further enhanced by our proposed method for determining the optimal step size. We evaluated our method and compared it to several state-of-the-art methods across various datasets. Our straightforward and effective approach consistently outperformed or achieved comparable performance to the state-of-the-art methods. Our code and datasets are available at https://github.com/zixinyang9109/BCF-FEM.
@@ I-B Contributions
In this paper, we introduce a 3D-3D non-rigid registration method based upon incorporating a modified FEM into the surface matching/data term. Most non-rigid registration methods incorporate the FEM model into the strain energy/regularization term. Only a few methods [16, 19] include the FEM in the surface matching data term, which, to our best knowledge, requires the manual identification and specification of boundary conditions.
Here, we alleviate the need to prescribe boundary conditions by using a modified stiffness matrix of the FEM, which is incorporated into the data term. We use a well-known numerical method (diagonal loading) to stabilize the stiffness matrix. As a result, the modified data term enables us to not need to manually identify and prescribe boundary conditions, and also to not need to use an additional strain energy term to be minimized.
@@ II-A Biomechanical Model
is a matrix used in each iteration of the solution of a nonlinear problem, and it changes with each iteration. For a linear elastic material, the tangent stiffness matrix is simply the stiffness matrix. A more detailed description of the construction of the FEM models can be in the following book
@@ II-B Stiffness Matrix and Boundary Conditions
To construct a FEM in linear elastic version as an example, where $\mathbf{f}=\mathbf{K}\mathbf{u}$, the values of $E$ and $\nu$ are used to construct the constant 3n x 3n stiffness matrix $\mathbf{K}$ [28, 29]. Given forces and ZBC, displacement $\mathbf{u}$ can be solved.
@@ III Methods
Our method employs a force-driven approach to deform the source model for optimal alignment with the target point cloud. Fig. 1 provides an overview of our registration method. The objective function comprises a surface matching term. The surface matching term utilizes a correspondence matrix generated by a closest point operator and a FEM. Minimization of the objective function is achieved through an accelerated gradient algorithm. During the force optimization phase, we initially determine the optimal step for calculating Nesterov’s accelerated gradient [30]. Subsequently, the gradient is employed to update forces. In each iteration, we first update the correspondence matrix via the closest point operator, then update the force through force optimization; finally, we update the displacements using the FEM. The FEM ensures that the displacement or deformation preserves the geometry of the preoperative model, maintaining the integrity of the node-to-node connections. After a set number of iterations, the predicted volumetric deformation enables the preoperative volume model to align with the intraoperative point cloud. In the following section, we will detail each of these steps.
@@ III-A Objective Function
where $\mathbf{C}\in\mathbb{R}^{3m\times 3n}$ is a soft correspondence matrix, $\mathbf{x}\in\mathbb{R}^{3n}$ are spatial location of the preoperative mesh nodes, $\mathbf{u}\in\mathbb{R}^{3n}$ are displacements associated with the nodes, and $\mathbf{y}\in\mathbb{R}^{3m}$ is the intraoperative point cloud. The correspondence matrix $\mathbf{C}$ maps the deformed model to points on the intraop surface will be detailed in Sec. III-B.
The liver biomechanical model can be incorporated into Eq. 4 by using the relationship between $\mathbf{u}$ and $\mathbf{f}$ given by Eq.2:
$J(\mathbf{f})=\frac{1}{2}\|\mathbf{C(u)}(\mathbf{x}+\mathbf{K}^{-1}\mathbf{f})-\mathbf{y}\|^{2}.$ (5)
The resulting data term now becomes a function of the forces applied to the surface of the liver. Our data term is a reformulation of the data term from Mestdagh et al. [16, 19] where we have explicitly incorporated the correspondence matrix and the dependence of $J(\mathbf{u})$ on $\mathbf{u}$.
@@ III-C Optimization
Nesterov’s accelerated gradient is given by Eq. 7 and Eq. 8. $\mathbf{p}^{k}$ carries momentum from the previous iteration by combining the force estimates of the current and previous iterations. Eq. 8 updates the forces, where $\nabla J(\mathbf{p}^{k})$ is the gradient of the data term (Eq. 5) with respect to $\mathbf{p}$, and $\alpha^{k}$ is the step size. $\nabla J(\mathbf{p}^{k})$ is given by
@@
Setting the derivative of Eq. 11 for $\alpha$ to zero, we can obtain the optimal step size:
@@ III-D Modifications for Unknown Zero Boundary Conditions and Force Locations
where $k_{ss}$ is the stiffness of the soft spring and $\mathbf{I}$ is the identity matrix. This is analogous to the work presented by Peterlik et al. [3], where a damping term was added to the stiffness matrix, which was then integrated into a strain/deformation energy term to regularize a data term. In our work, the modified stiffness matrix (Eq. 13) is integrated directly into the date term via the FEM. The added small stiffness increases the inertia of the liver model being deformed, preventing rigid motion without internal deformations. This approach enables us to allow the forces to be applied at any spatial location on the liver surface.
@@ IV-2 In vitro phantoms
Additionally, partial anterior surfaces were cropped to achieve surface visibility ratios ranging from approximately $18\%$ to $27\%$, calculated as the ratios of cropped surface areas to the full surface areas, as summarized in Table I.
@@ IV-3 Open-CAS dataset
The Open-CAS dataset contains three in silico phantoms and one in vitro phantom, as described in detailed in [5]. Deformations of the in silico phantoms are generated using a non-linear biomechanical model. The in vitro phantom featuring 6 fiducial markers was fabricated from silicone, and two intraoperative surfaces from a CT and stereo-endoscopic imaging are included.